# Eco-evolutionary dynamics of adapting pathogens and host immunity

### v2

@@ Abstract
To explore the behavior of our partial sweep model in a more general setting, we simulate evolutionary dynamics of a population under a Wright-Fisher model with expiring fitness dynamics. Simulations involve a population of $N$ viruses with a genotype where each position can be in one of two possible states $\sigma_{i}\in[0,1]$. Fitness effects $s_{i}$ are associated with mutations at each position, and the total fitness of a virus is given by $F=\sum_{i}\sigma_{i}s_{i}$. At each generation, viruses with a fitness $F$ expand by a factor $e^{F}$, and the next generation is constructed by sampling $N$ individuals from the previous one. Following Eq. 10, mutational effects $s_{i}$ decrease by an amount $\nu x_{i}\cdot s_{i}$, where $x_{i}$ is the frequency at which mutation $i$ is found in the population.
We simulate the emergence of adaptive variants in the following way. At a constant rate $\rho$, we pick one sequence position $i$ that has no polymorphism and set the fitness effect of the corresponding mutation to an initial value $s_{i}$, with an amplitude drawn from probability distribution $P_{s}$ and the sign chosen so as to favor mutations. In practice, we use an exponential distribution for $P_{s}$, meaning that the typical magnitude of initial fitness effects is described by only one parameter $s_{0}$. The corresponding distribution of partial sweep size is described in the SI. At the same time, we introduce the corresponding mutant in the population at a low frequency, picking its background genotype from a random existing strain. The behavior of the model is determined by (i) the distribution $P_{\beta}$ of partial sweeps size depending on $\nu/s_{0}$, and (ii) the ratio of the variant emergence rate and their growth rate $\rho/s_{0}$, which determines how often sweeps overlap and interact.
We use this simulation to address the question of predictability: given the state of the population at generations $0,1,\ldots,t$, can we predict its state at future times $t+1,\ldots$? Specifically, we ask whether we can predict the frequency $x(t+\Delta t)$ of a variant $A$, given it is at frequency $x$ at time $t$, as we did previously for influenza virus Barrat-Charlaix et al. (2021), see Fig. 3. The dynamics of isolated selective sweeps ($\rho/s_{0}\ll 1$, $\nu/s_{0}\ll 1$) should be perfectly predictable: after an initial stochastic phase when the variant is very rare, its frequency grows monotonically to fixation. This predictability decreases with increasing $\rho/s_{0}$ due to clonal interference (Strelkowa and Lässig, 2012), for example when an adaptive variant is outcompeted by an even more adaptive one. We also expect predictability to decrease with increasing $\nu/s_{0}$ since sweeps are then partial and their ultimate fixation determined by subsequent variants with dynamics that resemble a random walk.
@@
Figure 4: Simulations under the Wright-Fisher model with expiring fitness. A: Average frequency dynamics of immune escape mutations that are found to cross the frequency threshold $x^{*}=0.5$, for four different rates of fitness decay. If the growth advantage is lost rapidly (high $\nu/s_{0}$), the trajectories crossing $x^{*}$ have little inertia, while stable growth advantage (small $\nu/s_{0}$) leads to steadily increasing frequencies. B,C,D: Ultimate probability of $p_{fix}(x)$ of trajectories found crossing frequency threshold $x$. Each panel corresponds to a different rate of emergence of immune escape variants, with four rates of fitness decay per panel. Increased clonal interference $\rho/s_{0}$ and fitness decay $\nu/s_{0}$ both result in gradual loss of predictability. E: Time to most recent common ancestor $T_{MRCA}$ for the simulated population, as a function of the prediction obtained using the random walk $N_{e}=1/\rho\langle\beta^{2}\rangle$. Points correspond to different choices of parameters $s_{0}$
@@
In panels B, C and D of Figures 4, probability of fixations are shown for three values of $\rho/s_{0}$ and four values of $\nu/s_{0}$. Clonal interference increases when going from left to right among these panels (increasing $\rho/s_{0}$), while the intensity of fitness decay increases when going from blue to red curves (increasing $\nu$). Increasing either $\rho/s_{0}$ or $\nu/s_{0}$ reduces $p_{fix}$ towards the dashed diagonal corresponding to $p_{fix}=x^{*}$. However, as observed previously (Barrat-Charlaix et al., 2021), in the classic scenario with stable fitness effects $\nu/s_{0}=0$ considerable predictability remains even in cases of strong interference (blue curve in panel B).
@@
In panel E of Figure 4, the average time to the common ancestor of pairs of strains in the population is plotted as a function of $T_{2}$ predicted by the random walk model. Each point in the figure corresponds to one simulation of long duration with a given distribution of partial sweep size $P_{\beta}$ and a given $\rho$ setting $T_{2}$. We find a good agreement between the empirical time to MRCA and the estimation from the random walk, at least as long the probability of overlap between successive partial sweeps is small (indicated by shading, see SI). With increasing overlaps, coalescence slows down and diversity increases: points in darker shades of red tend to have larger time to MRCA than what is expected from the distribution of $\beta$. This is expected intuitively: if another adaptive variant emerges before the previous one has reached its final frequency, it has a lower probability to land on the same background and thus tends to be in competition with the first variant. This leads to a smaller effective $\beta$ which slows the dynamics.
@@
In any real-world scenario, there will a variety of mutations, including some mutations that perform complete selective sweeps, either because they escape immunity of a large fraction of the population ($M$ small), because they generate robust immunity against previous strains (“back-boost” (Fonville et al., 2014)), or because the increase the intrinsic transmissibility of the virus (for example reverting a previous escape mutation that had a deleterious effect on transmissibility). The degree to which partial sweeps matter will vary from virus to virus and will change over time. Recently emerged viruses circulate in a homogeneous immune landscape and adapt to the new host for some time, consistent with rapid and complete sweeps of variants in SARS-CoV-2. Similarly, influenza virus A/H1N1pdm, which emerged in humans in 2009, exhibited more consistent trajectory dynamics than A/H3N2 (Barrat-Charlaix et al., 2021).
@@ A.3 Equilibrium without the mutant
A.4 Realistic modeling of host’s immune state
@@
Figure S 1: Left: Dynamics of the frequency of the variant for the SIR model from Equations 23&24 using the invasion scenario from the main text. Two $2\times 2$ cross-immunity matrices are used, with off-diagonal parameters $f$ and $b$ chosen to give the same equilibrium. The gray line represents the equilibrium that would be obtained using the model of the main text. Right: Equilibrium frequency $\beta$ for this new SIR model ($y$-axis) versus the $\beta$ from the main text ($x$-axis). Each point corresponds to a given pair $(f,b)$.
A.5 Change in frequency when adding subsequent strains
@@
A.6 Case with intrinsic fitness effects
@@ A.7 Oscillations of the SI model
Appendix B Other
B.1 Differential equation with expiring fitness effects

### v3

@@ Abstract
To explore the behavior of our partial sweep model in a more general setting, we simulate evolutionary dynamics of a population under a Wright-Fisher model with expiring fitness dynamics. Simulations involve a population of $N$ viruses with a genotype where each position can be in one of two possible states $\sigma_{i}\in\{0,1\}$. Fitness effects $s_{i}$ are associated with mutations at each position, and the total fitness of a virus is given by $F=\sum_{i}\sigma_{i}s_{i}$. At each generation, viruses with a fitness $F$ expand by a factor $e^{F}$, and the next generation is constructed by sampling $N$ individuals from the previous one. Following Eq. 11, mutational effects $s_{i}$ decrease by an amount $\nu x_{i}\cdot s_{i}$, where $x_{i}$ is the frequency at which mutation $i$ is found in the population.
We simulate the emergence of adaptive variants in the following way. At a constant rate $\rho$, we pick one sequence position $i$ that has no polymorphism and set the fitness effect of the corresponding mutation to an initial value $s_{i}$, with an amplitude drawn from probability distribution $P_{s}$ and the sign chosen such that the mutation is adaptive. In practice, we use an exponential distribution $P_{s}\propto e^{-s/s_{0}}$, meaning that the typical magnitude of initial fitness effects is described by only one parameter $s_{0}$. The corresponding distribution of partial sweep size is described in the SI. At the same time, we introduce the corresponding mutant in the population at a low frequency, picking its background genotype from a random existing strain. The behavior of the model is determined by (i) the distribution $P_{\beta}$ of partial sweeps size depending on $\nu/s_{0}$, and (ii) the ratio of the variant emergence rate and their growth rate $\rho/s_{0}$, which determines how often sweeps overlap and interact. The probability of two sweeps overlapping is defined in Section (B.2) of the SI.
We use this simulation to address the question of predictability: given the state of the population at generations $0,1,\ldots,t$, can we predict its state at future times $t+1,\ldots$? Specifically, we ask whether we can predict the frequency $x(t+\Delta t)$ of a variant $A$, given it is at frequency $x$ at time $t$, as we did previously for influenza virus Barrat-Charlaix et al. (2021), see Fig. 3. The dynamics of isolated selective sweeps ($\rho/s_{0}\ll 1$, $\nu/s_{0}\ll 1$) should be perfectly predictable: after an initial stochastic phase when the variant is very rare, its frequency grows monotonically to fixation. This predictability decreases with increasing $\rho/s_{0}$ due to clonal interference (Schiffels et al., 2011; Strelkowa and Lässig, 2012), for example when an adaptive variant is outcompeted by an even more adaptive one. We also expect predictability to decrease with increasing $\nu/s_{0}$ since sweeps are then partial and their ultimate fixation determined by subsequent variants with dynamics that resemble a random walk.
@@
Figure 4: Simulations under the Wright-Fisher model with expiring fitness. A: Average frequency dynamics of immune escape mutations that are found to cross the frequency threshold $x^{*}=0.5$, for four different rates of fitness decay. If the growth advantage is lost rapidly (high $\nu/s_{0}$), the trajectories crossing $x^{*}$ have little inertia, while stable growth advantage (small $\nu/s_{0}$) leads to steadily increasing frequencies. B,C,D: Ultimate probability of $p_{fix}(x)$ of trajectories found crossing frequency threshold $x$. Each panel corresponds to a different rate of emergence of immune escape variants, with four rates of fitness decay per panel. Increased clonal interference $\rho/s_{0}$ and fitness decay $\nu/s_{0}$ both result in gradual loss of predictability. We use $s_{0}=0.03$. E: Time to most recent common ancestor $T_{MRCA}$ for the simulated population, as a function of the prediction obtained using the random walk $N_{e}=1/\rho\langle\beta^{2}\rangle$. Points correspond to different choices of parameters $\rho$ and $P_{\beta}$, and a darker color indicates a higher probability of overlap as computed in the SI.
@@
In panels B, C and D of Figures 4, probability of fixations are shown for three values of $\rho/s_{0}$ and four values of $\nu/s_{0}$. Clonal interference increases when going from left to right among these panels (increasing $\rho/s_{0}$), while the intensity of fitness decay increases when going from blue to red curves (increasing $\nu$). Increasing either $\rho/s_{0}$ or $\nu/s_{0}$ reduces $p_{fix}$ towards the dashed diagonal corresponding to $p_{fix}=x^{*}$. However, as observed previously (Barrat-Charlaix et al., 2021), in the classic scenario with stable fitness effects $\nu/s_{0}=0$ considerable predictability remains even in cases of strong interference (blue curve in panel D and Figure S3). The strong interference setting is explored in more detail in the SI up to values $\rho/s_{0}\simeq 30$, using similar simulations but without expiring fitness $\nu=0$. Figure S3 shows that even in these cases of strong interference, $p_{fix}$ remains significantly above the diagonal.
@@
In panel E of Figure 4, the average time to the common ancestor of pairs of strains in the population is plotted as a function of $T_{2}$ predicted by the random walk model. Each point in the figure corresponds to one simulation of long duration with a given distribution of partial sweep size $P_{\beta}$ and a given $\rho$ setting $T_{2}$, with darker color indicating a higher probability of overlap as computed in the SI. We find a good agreement between the empirical time to MRCA and the estimation from the random walk, at least as long the probability of overlap between successive partial sweeps is small (indicated by shading, see SI). With increasing overlaps, coalescence slows down and diversity increases: points in darker shades of red tend to have larger time to MRCA than what is expected from the distribution of $\beta$. This is expected intuitively: if another adaptive variant emerges before the previous one has reached its final frequency, it has a lower probability to land on the same background and thus tends to be in competition with the first variant. This leads to a smaller effective $\beta$ which slows the dynamics.
@@
In any real-world scenario, there will be a variety of mutations, including some mutations that perform complete selective sweeps, either because they escape immunity of a large fraction of the population ($M$ small), because they generate robust immunity against previous strains (“back-boost” (Fonville et al., 2014)), or because the increase the intrinsic transmissibility of the virus (for example reverting a previous escape mutation that had a deleterious effect on transmissibility). The degree to which partial sweeps matter will vary from virus to virus and will change over time. Recently emerged viruses circulate in a homogeneous immune landscape and adapt to the new host for some time, consistent with rapid and complete sweeps of variants in SARS-CoV-2. Similarly, influenza virus A/H1N1pdm, which emerged in humans in 2009, exhibited more consistent trajectory dynamics than A/H3N2 (Barrat-Charlaix et al., 2021).
@@ A.3 Equilibrium without the mutant
A.4 Equilibrium with $M$ immune groups
For $M$ immune groups and arbitrary cross-immunity matrices $\mathbf{K}_{i}$, the equilibrium frequency of the two strains is not easy to compute. However, it is possible to give an analytical expression in the regime of fast mixing $C_{ij}=M^{-1}$ and when the two strains differ immunologically for only one group, i.e. for matrices
$\mathbf{K}_{1}=\begin{pmatrix}1&b\\ f&1\end{pmatrix}\quad\text{and}\quad\mathbf{K}_{j}=\begin{pmatrix}1&1\\ 1&1\\ \end{pmatrix}\;\text{for}\;j>1.$
Note that this corresponds to the situation studied in the main text with fast mixing and $\varepsilon\rightarrow 0$. We show here that in this case, the equilibrium frequency for all immune groups is the same as the one obtained for only one immune group with matrix $\mathbf{K}_{1}$. In other words, the expression for $\beta$ in Eq. 6 is still valid.
To prove this, we assume the following form for the solution of the dynamical equations:
$\forall i\in\{1\ldots M\},\forall a,\;S_{i}^{a}=M\frac{\delta}{\alpha}\frac{I_{i}^{a}}{I^{a}}\quad\text{and}\quad\frac{I_{i}^{a}}{I_{i}}=\frac{I^{a}}{I}=\nu^{a},$ (24)
where the index $a$ runs over all strains (here wild-type and mutant), and where we have defined the infectious levels for group $i$, for strain $a$ and globally:
$I_{i}=\sum_{a}I_{i}^{a},\;I^{a}=\sum_{i}I_{i}^{a},\;I=\sum_{i,a}I_{i}^{a}.$
Note that the second equation in Eq. S24 means that the frequency $\nu^{q}$ of strain $a$ is the same accross all immune groups and consequently also globally.
We now show that injecting these expressions of $S$ and $I$ in the dynamical system and solving for $\nu^{a}$ gives the expected result. First, note that with this choice of $S_{i}^{a}$, the derivative of $I_{i}^{a}$ given by equation Eq. 1 immediatly vanishes. We thus concentrate on $\dot{S}_{i}^{a}$ given by Eq. 2. For any immune group $i$ and strain $a$, we have
$\dot{S}_{i}^{a}=0=-\delta\frac{I_{i}^{a}}{I^{a}}\sum_{b}K_{i}^{ab}I^{b}+\gamma\left(1-M\frac{\delta}{\alpha}\frac{I_{i}^{a}}{I^{a}}\right).$
where we have used $C_{ij}=1/M$ and $I^{b}=\sum_{j}I_{j}^{b}$ to remove the sum on immune groups. Multiplying this equation by $\nu^{a}=I^{a}/I$, we obtain
$-\delta I_{i}^{a}\sum_{b}K_{i}^{ab}\nu^{b}-\gamma\left(\nu^{a}-M\frac{\delta}{\alpha}\frac{I_{i}^{a}}{I}\right)=0.$
We now eliminate $I_{i}^{a}$ by using the expression $I_{i}^{a}=I_{i}\nu^{a}$:
$\delta I_{i}\nu^{a}\sum_{b}K_{i}^{ab}\nu^{b}-\gamma\left(1-M\frac{\delta}{\alpha}\frac{I_{i}}{I}\right)\nu^{a}$ $=0,$
$\delta I_{i}\sum_{b}K_{i}^{ab}\nu^{b}-\gamma\left(1-M\frac{\delta}{\alpha}\frac{I_{i}}{I}\right)$ $=0.$
Note that this last expression is true for all strains $a$. Considering any two strains $a$ and $b$, we can thus write
$\delta I_{i}\sum_{c}K_{i}^{ac}\nu^{c}-\gamma\left(1-M\frac{\delta}{\alpha}\frac{I_{i}}{I}\right)=0,$
$\delta I_{i}\sum_{c}K_{i}^{bc}\nu^{c}-\gamma\left(1-M\frac{\delta}{\alpha}\frac{I_{i}}{I}\right)=0,$
[…]
where the last expression is obtained by substracting the two previous ones. First, we see that for $i>1$ any frequency vector $\nu$ is a solution since $K_{i}^{ab}=1$ for all $a,b$. For $i=1$ and defining $\nu^{m}=\beta$ and $\nu^{wt}=1-\beta$ and using the expression for $\mathbf{K}_{1}$, we obtain
$(1-f)(1-\beta)+(b-1)\beta=0$
[…]
as claimed.
This result is not completely trivial and should be commented. In this setting, the mutant escapes immunity built by the wild-type for a fraction $1/M$ of the population, and yet it reaches the same frequency as in the case with one immune group. This can be rationalized as follows: for immune groups $i>1$, the cross-immunity matrix is such that the wild-type and mutant strains are completely equivalent. If immune group $1$ was not here, the mutant could thus equilibrate at any frequency between 00 and $1$. Since it is initially introduced at very low frequency, it would remain marginal in immune groups $i>1$. However, since its “natural” equilibrium frequency in group $i=1$ is $\beta$ and since the groups are connected, equilibrium is reached when the mutant reaches frequency $\beta$ in all groups.
Note that if we take the situation of the main text with
$\mathbf{K}_{i}=\begin{pmatrix}1&\varepsilon\\ \varepsilon&1\end{pmatrix}$
and $\varepsilon>0$, the expressions above do not hold. However, if $\varepsilon\ll 1$, the perturbation is small and we expect an equilibrium frequency close to $\beta$, which the case in Figure 1.
A.5 Realistic modeling of host’s immune state
@@
Figure S 1: Left: Dynamics of the frequency of the variant for the SIR model from Equations 25&26 using the invasion scenario from the main text. Two $2\times 2$ cross-immunity matrices are used, with off-diagonal parameters $f$ and $b$ chosen to give the same equilibrium. The gray line represents the equilibrium that would be obtained using the model of the main text. Right: Equilibrium frequency $\beta$ for this new SIR model ($y$-axis) versus the $\beta$ from the main text ($x$-axis). Each point corresponds to a given pair $(f,b)$.
A.6 Change in frequency when adding subsequent strains
@@
A.7 Case with intrinsic fitness effects
@@ A.8 Oscillations of the SI model
A.9 Link between parameters of the SIR and expiring fitness models
The effective expiring fitness model used in the second part of this work is characterized by the system of differential equations
$\dot{x}=sx(1-x)\quad\text{and}\quad\dot{s}=-\nu xs,$
where $x$ is the frequency of the mutant strain. Here, we try to express the dynamics of the SIR model in this form to find a link between its parameters and the quantities $s$ and $\nu$.
We first focus on the case with two strains and one immune group. The frequency of the mutant is $x=I^{m}/(I^{m}+I^{wt})$. Using the logit function $\psi(x)=\log(x/(1-x))$ and the dynamical equations of the SIR, we find
$\frac{\text{d}}{\text{d}t}\psi(x)=\alpha(S^{m}-S^{wt}),$ (37)
which allows us to define the fitness in the SIR case: $s=\alpha(S^{m}-S^{wt})$. At the beginning of the invasion, the initial growth rate is readily computed:
$s(t=0)=\frac{(1-f)(\alpha-\delta)}{\delta+f(\alpha-\delta)}\delta,$ (38)
which is the same as the initial growth rate of $I^{m}$. Note that if $f=1$, the initial growth rate is 00.
We then compute the time derivative of $s$ early in the invasion, when $I^{wt},S^{wt}$ and $S^{m}$ are close to their equilibrium values. In this case, a straightforward calculation gives
$\begin{split}\dot{s}=&-\alpha(I^{wt}+I^{m})\cdot\alpha(S^{m}-bS^{wt})\cdot x\\ \simeq&-\alpha(I^{wt}+I^{m})xs,\end{split}$ (39)
where the approximation is valid if $1-b\ll 1$. This would give an expiry rate of fitness $\nu=\alpha(I^{wt}+I^{m})$ in the case of the SIR model.
These results can also be obtained in the case of immune groups. We then have the following expressions for $s$ and $\nu$ at $t\simeq 0$:
$\begin{split}s=&\frac{\alpha}{M}\sum_{i=1}^{M}(S_{i}^{m}-S_{i}^{wt})\\ \dot{s}=&-\alpha(I^{wt}+I^{m})x\sum_{i=1}^{M}(S_{i}^{m}-b_{i}S_{i}^{wt})\\ =&-\alpha(I^{wt}+I^{m})xs,\end{split}$ (40)
where the second expression is again valid if $1-b_{i}\ll 1$.
Another question is that of the link between cross-immunity parameters $b$ and $f$ and the distribution of partial sweep sizes $\beta$. The relation between cross-immunity and $\beta$ given by Eq. 6 shows that the distribution of partial sweep size depends on the distribution of both $b$ and $f$. As we do not have a prior on how $b$ and $f$ should be distributed, we explore the case where $1-f$ and $1-b$ are exponentially distributed. In other words, we define $\epsilon_{f}=1-f$ and $\epsilon_{b}=1-b$, with the following distributions
$P(\epsilon_{f})\propto e^{-\epsilon_{f}/\lambda}\quad\text{and}\quad P(\epsilon_{b})\propto e^{-\epsilon_{b}/\mu}.$
The expression of the partial sweep size becomes $beta=\epsilon_{f}/(\epsilon_{b}+\epsilon_{f})$. Note that both $\epsilon$ should remain smaller than $1$, which is not guaranteed with exponential distributions. However, this is not problematic if $\mu$ and $\lambda$ take small enough values.
These assumptions allow us to compute the distribution of $beta$:
$P(\beta)=\frac{\mu/\lambda}{(\frac{\mu}{\lambda}\beta+(1-\beta))^{2}}$
with support on the interval $0\leq\beta\leq 1$. Figure S2 shows the various shapes that $P(\beta)$ then takes for different values of the $\mu/\lambda$ parameter. Note that if $\mu>\lambda$, $b$ tends to be higher than $f$ and $\beta$ is biased towards one. If $\mu=\lambda$, $P(\beta)$ becomes uniform on the $[0,1]$ range.
Figure S 2: Distribution of partial sweep size $\beta$ if $1-f$ and $1-b$ are exponentially distributed. Left: Probability distribution function $P(\beta)$ for various values of $\mu/\lambda$. Right: Mean and standard deviation of $\beta$ as a function of $\mu/\lambda$.
The exponential distribution away from one of $f$ and $b$ is a reasonable assumption, and allows analytical derivation of $P(\beta)$. Of course, any other distribution of $f$ and $b$ could be considered. For this reason, we choose to use a Beta distribution for $P_{\beta}$ in the analysis of the main text, as it can accomodate various shapes.
Appendix B Expiring fitness model and random walk
B.1 Clonal interference
In non recombining genomes with a large mutation rate, the appearance of many adaptive mutations in close succession leads to clonal interference. In this regime, beneficial mutations present on different background compete for fixation, and the success of a mutation does not depend only on its fitness effect but also on the global state of the population. For this reason, clonal interference causes a decrease in predictability: dynamics are not deterministic but rather depend on the precise structure of which mutation appeared on which background. For instance, a beneficial mutation that increases in frequency can be outcompeted by a fitter one before it has the time to fix, making the extrapolation of frequency trajectories difficult.
We conduct simulations to quantify how much predictability decreases because of clonal interference, based on the ones of the main text. We study a large population of $N=10^{5}$ genomes of length $L$, where each genome position $i$ can be in either of two states $\sigma_{i}\in\{0,1\}$. Fitness effects $s_{i}\in\mathbb{R}$ are associated to each position, and the fitness of an individual is $F=\sum_{i}s_{i}\sigma_{i}$. To simulate the adaptation of the population, we proceed in the following way: at a constant rate $\rho$, we pick a position $i$ that is non polymorphic and set the fitness effect $s_{i}$ by sampling its magnitude from distribution $P_{s}$ and choosing its sign in a way that favors mutations (positive if $\sigma_{i}=0$ is more frequent, negative otherwise). At the same time, the corresponding mutation ($\sigma_{i}=0$ if $s_{i}<0$ and inversely) is introduced in the population at a small frequency $\delta f=0.01$. We choose $P_{s}$ to be an exponential distribution with scale $s_{0}=0.02$, in agreement with findings on the distribution of fitness effects in Schiffels et al. [2011], Rice et al. [2015].
There is no fitness decay in this simulation, and the parameters $N$, $s_{0}$ and $\delta f$ have numerical values such that neutral drift is mostly irrelevant to the dynamics of mutations. For example, without accounting for interference, the probability of fixation of a mutation of effect $s_{0}$ when it is introduced at frequency $\delta f$ is $p_{fix}(\delta f)\simeq 1-e^{-40}$. As a consequence, the two parameters governing the dynamics are the scale of fitness effects $s_{0}$ and the rate of introduction of beneficial mutations $\rho$. The ratio of these two quantities represents the amount of clonal interference: at low $\rho/s_{0}$, mutations are mostly independent, while at high $\rho/s_{0}$ they strongly interfere.
We measure the probability of fixation $p_{fix}(x)$ of mutations found in a frequency bin $[x-\delta x,x+\delta x]$ over a long simulation. We only consider mutations with increasing frequency, meaning that their frequency was below $x$ at all times and at some point was measured in the frequency bin. Figure S3 shows $p_{fix}(x)$ as a function of $x$ for different values of $\rho/s_{0}$. According to intuition, a low clonal interference value leads to easily predictable fixations: whatever the frequency $x$ at which it is observed, a mutation that is increasing in frequency fixes with a very high probability. Increasing clonal interference clearly makes dynamics less predictable and closer to neutrality, with $p_{fix}$ approaching the diagonal line. However, even in a regime of strong interference, e.g. $\rho/s_{0}=32$, deviations from neutrality remain very clear.
Figure S 3: Probability of fixation of mutations $p_{fix}(x)$ of mutation frequency trajectories found crossing the frequency threshold $x$. Fitness effects are exponentially distributed with fixed scale $s_{0}=0.03$. The blue to red gradient in colors corresponds to the increasing rate $\rho$ at which mutations are introduced. Strong clonal interference regime is obtain when $\rho/s_{0}>1$, in which case good mutations are introduced in close succession and compete for fixation. At low $\rho/s_{0}$, trajectories are very predictible and an increasing trajectory almost certainly fixes. Even for for the highest $\rho/s_{0}$, $p_{fix}(x)$ remains significantly larger than $x$ and dynamics are visibly not neutral.
B.2 Expiring fitness effects: sweep size and probability of overlap